OpenStudy (anonymous):

How much is 17% of 30? 0.51 5.1 5.7 51

OpenStudy (anonymous):

5.1

OpenStudy (anonymous):

17/100 * 30 = 17/10 * 3 = 1.7*3 = 5.1